# Daria Semochkina School of Mathematical Sciences

### v1

@@ Abstract
Background Statisticians evaluating the impact of policy interventions such as screening or vaccination will need to make use of mathematical and computational models of disease progression and spread. Calibration is the process of identifying the parameters of these models, with a Bayesian framework providing a natural way in which to do this in a probabilistic fashion. Markov Chain Monte Carlo (MCMC) is one of a number of computational tools that is useful in carrying out this calibration. Objective In the context of complex models in particular, a key problem that arises is one of non-identifiability. In this setting, one approach which can be used is to consider and ensure that appropriately informative priors are specified on the joint parameter space. We give examples of how this arises and may be addressed in practice. Methods Using a basic SIS model the calibration process and the associated challenge of non-identifiability is discussed. How this problem arises in the context of a larger model for HPV and cervical cancer is also illustrated. Results The conditions which allow the problem of non-identifiability to be resolved are demonstrated for the SIS model. For the larger HPV model, how this impacts on the calibration process is also discussed.
@@ 1 Introduction
As statisticians, we are often involved as part of a larger team of applied mathematicians, computer modellers and epidemiologists when fitting and using disease models such as those described in this article.
The use of mathematical models to predict likely outcomes of different medical interventions, such as vaccination, screening programs and treatments to tackle a disease can be extremely helpful. Computer-based models of disease spread and progression are used for medical decision making, economic and policy evaluation in disease evaluation and control. In the case of health systems, decision makers are often doctors or administrators, making decisions on behalf of patients, current and future users and of the health system.
The key issue examined in this paper where the input of the statistician is important is the area of model calibration. We examine an approach to calibration, how the stochastic samples from the parameter sets are summarised and how to deal with a key challenging aspect, namely that of non-identifiability.
In healthcare applications, model calibration is the process of identifying a set or sets of parameters so that the mathematical model yields predictions for population characteristics (e.g. mortality, number of cases of a disease, the severity of illness) that match what is actually observed. As there is a great demand for model calibration techniques, there exist numerous methods and approaches [1].
In model calibration, so-called ‘calibration targets’ are usually chosen to correspond to some data that are observed directly. The aim of Bayesian model calibration is to find a distribution for unobservable parameters (model inputs), conditional on available data (calibration targets). This is often referred to as an inverse problem. In some situations, it may be feasible and sufficient to obtain a set of parameters using a best-fit approach or a pure optimisation technique. However, these approaches underestimate the uncertainty associated with any forecasts that may be made with the model by only providing point estimates. It is therefore desirable that this uncertainty in the distribution of the parameters obtained through calibration be quantified somehow. In addition, some frequently adopted approaches such as grid search produce worse fits of the model to the observed data when using the same computational resources[2].
The Bayesian framework is a natural one to use for solving inverse problems in general and consequently perfect for model calibration. The advantages of the Bayesian approach include conditional sampling and uncertainty quantification. The main computational complexity in Bayesian inversion is in sampling from a posterior distribution. The distribution of interest is often high-dimensional and does not have a standard form. Markov chain Monte Carlo (MCMC) methods are powerful tools for solving inverse problems and obtaining a sample from the posterior distribution of interest[2].
We focus this paper on solving inverse problems[3], illustrating it with a human papillomavirus (HPV) model calibration example. While other approaches exist for sampling from Bayesian posterior distributions, here we use MCMC algorithms, in particular the Metropolis-Hastings algorithm. This approach and this example is useful from the point of view of illustrating generic calibration problems, as well as being a realistic model for a topic of substantive interest.
The Bayesian approach involves obtaining a probability distribution for the parameters of interest, $\theta$, having seen some data, $\mathscr{D}$. In this case, the data are the ‘target’ observations, e.g. prevalence of disease at a particular time point. It is a necessary feature of the paradigm that ‘prior’ distributions for the parameters of interest are formulated, which encapsulate our understanding of the parameter space without directly incorporating evidence from the data.
Once a prior has been specified, and targets have been obtained, Bayes Theorem allows us to obtain an expression for the posterior distribution.
$\mathscr{P}(\theta|\mathscr{D})\propto\mathscr{P}(\mathscr{D}|\theta)\cdot\mathscr{P}(\theta),$
where $\mathscr{P(\theta)}$ is the prior distribution and $\mathscr{P}(\mathscr{D}|\theta)$ is the likelihood function. All of the information relevant to $\theta$ contained in the targets (here called $\mathscr{D}$) is contained in the likelihood function. Other information about our parameters can be summarised in the prior. Where there is no additional information the prior is often specified as a uniform distribution, in which case the posterior has the same shape as the likelihood. Of particular importance here is that information which is not contained in the targets, that is relevant to the inference may be incorporated through the prior. As will be seen later, a reconsideration about what external information about $\theta$ exists may be useful in addressing the question of non-identifiability.
In the next section, we discuss the methodology of calibration in general, introduce the Metropolis-Hastings algorithm and discuss potential limitations. Some approaches to resolve non-identifiability is discussed in Section 3. We then proceed to examine this in the context of a simple SIS disease transmission model in Section 4. We illustrate the role of the prior here in inducing identifiability during the calibration process. In Section 5, we present the HPV model and the Ireland-specific data to which the model is calibrated. In Section 6, we present simulation results with a few modifications of the Metropolis-Hastings algorithm to improve the performance, including an examination of the effects of having an informative prior distribution. Finally, Section 7 provides some discussion and conclusions.
2 Model calibration using the Metropolis-Hastings algorithm
If we formulate the model calibration problem in a probabilistic way, the output of model calibration will be a sample from the parameters’ a posteriori probability distribution. Assuming that each parameter combination produces a unique forward model causes the outcome of model calibration to not be a unique model but rather a collection of models. A single model could be the result of techniques that produce a single best-fitting point estimate in the parameter space. These models can be equally consistent with the observed data. This collection should agree with both the a priori information—incorporated in the model through prior distributions—and the observed data, accurately incorporating the uncertainty about observations. The Metropolis-Hastings algorithm searches through the joint parameter space until the chain converges in distribution around sets of values (the stationary distribution of the Markov chain is reached), which may then be used for inference or decision making.
@@ 2 Model calibration using the Metropolis-Hastings algorithm
The Metropolis-Hastings algorithm (M-H) is an MCMC algorithm for obtaining a random sample from a probability distribution $f$, for which direct sampling is difficult. The M-H algorithm has been tremendously successful in modern computational statistics because it enables sampling from complex distributions to obtain independent samples directly from such distributions. The algorithm is based on proposing values, sampled from some proposal distribution. These values are then accepted or rejected with a certain probability of acceptance, based on how likely those values are for some distribution of interest $f$. See Appendix A.1 for details.
@@ 2.1 Challenges
A number of challenges exist when calibrating models using limited data. These include difficulties related to the high dimensionality of the parameter space, multi-modality of the likelihood function and the question of identifiability.
When a very large number of parameters are present (i.e. high dimensionality) this increases the complexity and the amount of computation required regardless of the numerical algorithm used.
Where multiple modes (local maxima) are present in the likelihood, algorithms risk ‘getting stuck’ at local optima.
The challenge of identifiability[4] arises where the data are consistent with at least two (often a continuum of) different sets of parameter values. For example in the case of disease progression, the data could be consistent with situations where the rate of progression is fast and the rate of recovery is fast but equally, it could be the situation where the rate of recovery is slow with a slow rate of progression.
Some models (e.g. micro-simulation or agent based models) are computationally intensive, and so it is necessary to ensure the sampling algorithm is efficient
3 Approach to resolving non-identifiability
@@
In practice, non-identifiability can be diagnosed when there is a lack of convergence of chains, when the posterior distributions of parameters are flat and broad, and when there is cross-correlation between sampled values.
@@
In order to resolve non-identifiability, the suggested approach is to include information about some of the variables where this is obtainable. This can be done by carefully considering whether additional information has been collected or is available from other sources. For example duration of illness may not be available locally, but a prior may be obtained by considering international evidence that is emerging. In algorithmic terms the process of calibration would proceed as follows:
@@
  • Identify targets and populate these with collected data.
@@
  • Run MCMC calibration process.
@@ 4 Example: SIS model
To illustrate this further we provide an example in line with the model introduced in Figure 1(a). We consider a simple susceptible-infected-susceptible (SIS) compartmental model. The transitions between the two states are described by
@@
where $\beta=c\cdot p$ and $\gamma=1/d$ with $c$ being the contact rate, $p$ the transition probability and $d$ the infectious period. Most of the population starts off in the susceptible state with $S=0.99$ and $I=0.01$. We treat this situation as a population where there has initially been disease introduced, and at some later stage, a stable situation has been observed where 1200 of 2000 individuals are identified as infected. Thus, the likelihood may be described as having a similar shape to normal distributions $N(0.4,0.01)$ and $N(0.6,0.01)$, and these targets are then used in the algorithm. Implicit in this is that an improper uniform prior is used for p and c.
@@
After running a simple Metropolis-Hastings algorithm for $c,\ p$ and $d$ with those normal distributions as our targets we plot the resulting $\beta$ and $\gamma$ values in Figure 1(b). It is noticeable that those are almost perfectly correlated. In panel (b), of the figure, the impact of specifying a prior for one variable is shown. Even though no information about the second variable is added, conditioning on a narrower prior distribution for $\gamma$ restricts the joint posterior space to a narrower range, thus impacting the marginal posterior distribution for $\beta$ which is induced due to the correlation between the parameters. Additionally, because $\beta=c\cdot p$, $c$ and $p$ are perfectly correlated. This leads to the chains for $c,\ p$ and $d$ to not converge due to the complex links between their values and the observational equivalence of the resulting targets.
(a) Schematic of the system.
(b) Transition parameter values fitting the observed targets well.
Figure 1: Simple example of an SIS model. Parameters $\beta$ and $\gamma$ are transition parameters. Note the density in panel (b) represents an informative prior, with arrows illustrating how it impacts upon the joint posterior and hence the marginal posterior of $\gamma$. The annotated density on the x axis for $\beta$ illustrates how the information in that parameter restricts the joint posterior, and thus provides indirect information on $\gamma$, the marginal distribution of which is now affected.
@@
Figure 2: Resulting chains of running the Metropolis-Hastings algorithm with improper uniform priors and with informative priors.
As identified by Alarid-Escudero et al. [4], the main problem in models of this type is the lack of identifiability in the parameter space, given the information available. One solution is to obtain additional data, or an additional target for which data may be available.

### v2

@@ Abstract
Disease models are used to examine the likely impact of therapies, interventions and public policy changes. Ensuring that these are well calibrated on the basis of available data and that the uncertainty in their projections is properly quantified is an important part of the process. The question of non-identifiability poses a challenge to disease model calibration where multiple parameter sets generate identical model outputs.
For statisticians evaluating the impact of policy interventions such as screening or vaccination, this is a critical issue. This study explores the use of the Bayesian framework to provide a natural way to calibrate models and address non-identifiability in a probabilistic fashion in the context of disease modelling. We present Bayesian approaches for incorporating expert knowledge and external data to ensure that appropriately informative priors are specified on the joint parameter space. These approaches are applied to two common disease models: a basic Susceptible-Infected-Susceptible (SIS) model and a much more complex agent-based model which has previously been used to address public policy questions in HPV and cervical cancer. The conditions which allow the problem of non-identifiability to be resolved are demonstrated for the SIS model. For the larger HPV model an overview of the findings is presented, but of key importance is a discussion on how the non-identifiability impacts the calibration process. Through case studies, we demonstrate how informative priors can help resolve non-identifiability and improve model inference. We also discuss how sensitivity analysis can be used to assess the impact of prior specifications on model results. Overall, this work provides an important tutorial for researchers interested in applying Bayesian methods to calibrate models and handle non-identifiability in disease models.
@@ 1 Introduction
Disease modelling plays a crucial role in epidemiology by providing insights into disease transmission dynamics and informing public health interventions. Mathematical models capture these dynamics through a set of parameters, such as rates related to transmission, recovery, and mortality. Such parameters quantify key features of the disease. Estimating those parameters accurately is essential for reliable model predictions. The process of estimating model parameters is termed calibration. This work was motivated by the authors’ previous experience dealing with challenges which arose while calibrating disease models for the purposes of informing policy on screening programmes[1], vaccination strategies[2] and while involved in modelling during the COVID pandemic[3].
Bayesian inference offers a powerful framework for parameter estimation in disease models. It is often desirable that the uncertainty associated with any forecasts is quantified in the distribution of the parameters obtained through calibration. The Bayesian framework is a natural one to use for solving inverse problems such as model calibration. The goal of Bayesian inference is to sample from a probability distribution which is most compatible with a combination of data and a prior distribution describing the parameter space. The distribution of interest is often high-dimensional and does not have a standard form. Markov chain Monte Carlo (MCMC) methods are powerful tools for solving inverse problems and obtaining a sample from the posterior distribution of interest[1]. We examine the Bayesian approach to calibration, how the stochastic samples from the parameter sets are summarised and how to deal with a key challenging aspect, namely that of non-identifiability.
A model is said to be non-identifiable if two or more distinct sets of parameter values can produce the same observed data. That is, we cannot uniquely determine the true parameter values from the available data. For example, in a simple linear regression model, if two predictor variables are perfectly correlated, it becomes impossible to isolate the individual effects of each variable on the response. The issue of identifiability can occur in basic mixture models (such as a simple two-component mixture) and has been seen by the current authors in other situations, such as that of label switching in latent variable models [4]. This can lead to difficulty in interpreting posterior distributions and drawing reliable inferences about the model parameters. In the context of disease modelling, non-identifiability can occur for various reasons, such as limitations in the available data or inherent structural features of the model[5, 6, 7, 8, 9].
This study addresses the problem of non-identifiability in Bayesian disease model calibration. We explore the use of informative prior information, derived from expert knowledge or external data, to resolve non-identifiability issues. We focus on two common disease models: the Susceptible-Infected-Susceptible (SIS) model and the agent-based Susceptible-Infected-Recovered-Deceased (SIRD) HPV model with probabilistic transition rates. We will demonstrate how informative priors can help to improve the precision of posterior estimates. Additionally, we discuss the importance of sensitivity analysis to evaluate the robustness of model results to variations in prior specifications.
2 Disease Models and Bayesian Model Calibration Basics
This section introduces the fundamental concepts of disease modelling and the Bayesian approach to model calibration. We explore how Bayesian inference provides a powerful framework for parameter estimation in disease models. We outline the challenges that arise with non-identifiability in model calibration and common causes of non-identifiability. Several recommendations are provided to tackle non-identifiability.
2.1 Compartmental Models for Disease Dynamics
Disease models often use a compartmental approach to describe the progression of infections within a population. These models categorise individuals into compartments based on their disease status and define transitions between these states through mathematical equations. This framework has been instrumental in understanding the dynamics of disease transmission, evaluating control measures, and guiding public health policy[10].
One of the earliest examples of compartmental models for disease dynamics is the SIR model, introduced by Kermack and McKendrick[11] in 1927, which includes states of susceptible (S) and infectious (I) as well as an additional compartment: recovered (R). In the SIR model, individuals transition from susceptible to infected and then to recovered, where they gain immunity. The SIR model underpins many critical epidemiological concepts, such as the basic reproduction number, $R_{0}$, and the herd immunity threshold. This model is widely used for studying outbreaks of infectious diseases like measles and influenza, where immunity plays a significant role in the population’s response to an epidemic.
Another foundational example is the SIS model[12], where individuals are grouped into two compartments: susceptible (S) and infected (I) (see Figure 1(a)). Susceptible individuals can contract the disease through contact with infected individuals at a rate defined by the transmission parameter $\beta$. Infected individuals recover at a rate $\gamma$ but lose immunity and return to the susceptible state. This structure is ideal for diseases where immunity is not long-lasting, such as certain bacterial infections.
For diseases with an incubation period, the SEIR model introduces an exposed (E) compartment, representing individuals who are infected but not yet infectious[13]. This extension captures the dynamics of diseases like COVID-19[14, 15, 16] or influenza[17, 18]. Individuals move from the susceptible state to the exposed state upon infection at a rate $\beta$, then progress to the infected state after an incubation period at rate $\sigma$, and finally recover at rate $\gamma$.
The work of Kermack and McKendrick also extended to include endemic diseases by incorporating demographic factors such as births, deaths, and partial immunity[19, 20]. These models provide a framework for analysing diseases that persist in populations over long periods, such as malaria or tuberculosis, and are capable of capturing the effects of interventions like vaccination.
Compartmental models remain indispensable for understanding disease dynamics, predicting epidemic trajectories, and guiding public health interventions. They have been applied in diverse contexts, from classical studies of measles outbreaks to the global response to the COVID-19 pandemic. For example, the SEIR model was crucial in forecasting the impact of social distancing and vaccination strategies during the early stages of COVID-19[21, 22]. Similarly, the foundational SIR model continues to inform vaccine allocation and predict outbreak sizes. These applications highlight the flexibility and enduring relevance of compartmental models in addressing emerging and existing infectious diseases
2.2 Parameter Estimation and Key Concepts in Bayesian Inference
Estimating parameters (e.g., transmission rate, recovery rate) of a disease model allows us to understand and make predictions about the disease dynamics. Bayesian inference helps us quantify our knowledge about these parameters in terms of a joint probability distribution. In general, Bayesian inference provides a statistical framework for reasoning under uncertainty. It has the following key constructs:
  • Prior Distribution: Represents our beliefs or knowledge about the parameter values without consideration of directly observed data. Priors can be informative (based on expert knowledge or previous studies) or non-informative (weak priors with minimal influence). Information in the prior is separate from that obtained during a study or experiment used for model fitting or calibration.
  • Likelihood Function: The likelihood is the function through which new data inform parameter estimates. It quantifies the probability of observing the data given specific parameter values. The higher the likelihood, the better the data fits the model with those parameter values.
  • Posterior Distribution: Combines the prior information with the information from the newly observed data. It represents the updated beliefs about the parameter values after considering both the prior and the observed data. The posterior distribution is fully specified by the combination of prior and likelihood.
In the context of disease modelling the prior distribution could represent expert knowledge about the typical range of values taken by the transmission rate of a disease. The likelihood function would quantify how well a specific transmission rate value fits the observed disease prevalence data. The posterior distribution would give us a more precise estimate of the transmission rate, taking into account both prior knowledge and the data.
2.3 Implementation of Bayesian Model Calibration
There are various software implementations and computational methods for implementing Bayesian model calibration [23, 24] with our implementation here using Markov Chain Monte Carlo (MCMC) methods[25, 26, 27, 28][29, 1, 30]. These methods involve generating a large number of sets of parameter values from the posterior distribution which can be used to draw inferences about the model parameters.
@@ 2.3 Implementation of Bayesian Model Calibration
The Metropolis-Hastings algorithm (M-H)[31, 32] is an MCMC algorithm for obtaining a sample from a probability distribution $f$, for which direct sampling is difficult or impossible. The M-H algorithm enables sampling from complex distributions to obtain (almost) independent samples directly from such distributions[33, 27]. The algorithm is based on proposing values, sampled from some proposal distribution. These values are then accepted or rejected with a certain probability of acceptance, based on how likely those values are for some distribution of interest $f$, generating a Markov chain of parameter values that eventually converges to the target distribution. See Appendix A for details.
The steps in the Metropolis-Hastings Algorithm are outlined as follows. Step 1 (Initialisation): Choose a starting value for the parameter vector $\boldsymbol{x}$. Step 2 (Proposal Generation): Propose a new parameter vector based on a proposal distribution (e.g., a random walk, a multivariate normal distribution, etc.). Step 3 (Acceptance Probability): Calculate the acceptance probability based on the ratio of the posterior densities at the proposed and current parameter values. Note that because the algorithm uses a ratio, a normalising constant is not required[32], which is the usual obstacle in acquiring the posterior distribution. Step 4 (Acceptance or Rejection): Randomly decide whether to accept or reject the proposed value based on the acceptance probability. If accepted, the proposed value becomes the new state of the chain; otherwise, the current state is retained. Iteration: Repeat steps 2-4 for a specified number of iterations.
Some practical considerations are worth reviewing here.
  • Firstly, the choice of proposal distribution can significantly impact the efficiency of the algorithm. A well-tuned proposal can improve mixing and reduce autocorrelation of the generated sample.
  • Secondly, it is essential to assess whether the MCMC chain has converged to the target distribution. Visual diagnostics of trace plots and tools like the Gelman-Rubin statistic can be used for this purpose[34].
  • Thirdly, the initial samples from the MCMC chain may not be representative of the target distribution. A burn-in period is often used to discard these early samples.
  • Finally, to reduce the autocorrelation of the samples, the chain can be thinned by selecting every $n$-th point.
Several statistical software packages (e.g., R[35], Stan[36], WinBUGS[37]) provide implementations of the Metropolis-Hastings algorithm and other MCMC methods. Regardless of the software used, the considerations about non-identifiability arise, and the solution using our proposed approach applies. In this tutorial, we will focus on implementing the Metropolis-Hastings algorithm in particular and addressing these practical considerations. We will demonstrate that by doing this, researchers can obtain reliable parameter estimates for their disease models, even in the presence of non-identifiability.
@@ 2.4 Challenges
A number of challenges exist when calibrating models using limited data. These include difficulties related to the high dimensionality of the parameter space, multi-modality of the likelihood function, and, more importantly, the question of identifiability for model calibration. High-dimensional parameter spaces: When models involve numerous parameters, the computational burden of exploring the posterior distribution can be substantial. Multi-modality: The likelihood function may have multiple local maxima, leading to difficulties in convergence. Computational intensity of models: Models like micro-simulation or agent-based models can be computationally expensive to simulate and it is necessary to ensure the sampling algorithm is efficient.
The challenge we will focus our attention on in this tutorial is that of identifiability[5]. As mentioned before, non-identifiability arises where the data are consistent with at least two (often a continuum of) different sets of parameter values. For example in the case of disease progression, the data could be consistent with situations where the rate of progression is fast and the rate of recovery is also fast but equally, it could be the situation where the rate of recovery is slow with a slow rate of progression. In this situation, commonly collected data such as that on the prevalence of a disease gives only the total number of cases of this disease in a given population at a specific time. This data will be consistent with individuals progressing and recovering quickly as well as being consistent with individuals both progressing and recovering slowly. No information in the prevalence data can distinguish between those scenarios. In this situation, a good solution to overcoming the question of identifiability of the progression rate would be to elicit prior information about the recovery rate either from experts or through direct observation of cases.
2.5 Non-Identifiability in Disease Models
Non-identifiability arises when multiple parameter sets can generate the same model output. In the context of disease models, different model parameter combinations can produce indistinguishable disease dynamics patterns. There are several examples of non-identifiability. Simultaneous changes in incidence and recovery rates (i.e. increasing both the incidence rate and the recovery rate) can lead to similar epidemic curves, making it difficult to distinguish between these two effects. Models with hidden states, such as latent infections or asymptomatic cases, can introduce non-identifiability if the data do not directly observe these states. Structural non-identifiability occurs when the model structure itself is inherently non-identifiable, regardless of the data. For example, certain compartmental models may have symmetries that lead to non-identifiability.
There are some major implications of non-identifiability[5]. Non-identifiability can lead to wide credible intervals or flat posterior distributions, making it difficult to draw precise inferences about model parameters. If non-identifiability is not addressed, it can lead to incorrect model predictions and misleading conclusions. Non-identifiable models can be difficult to calibrate using standard methods, as the likelihood surface may have multiple modes or flat regions.
Several methods can be used to detect non-identifiability.
  • Profiling the likelihood function over different parameter values could help identify regions of the parameter space where the likelihood is relatively flat.
  • The Fisher information matrix can be used to assess the curvature of the likelihood surface and identify regions of non-identifiability. However, this is not an option for non-parametric models which are the most used in disease modelling.
  • Introducing additional latent variables can sometimes help to resolve non-identifiability issues.
  • It was also demonstrated that using multiple observed responses that all depend on some common set of calibration parameters[9] could in some cases improve identifiability.
Several approaches can be applied to addressing non-identifiability. Incorporating strong prior information can help to break the degeneracy and improve identifiability. In some cases, it may be possible to reformulate the model to address non-identifiability. For example, introducing additional constraints or assumptions can help to narrow down the parameter space. Adding additional data or measurements can sometimes help to resolve non-identifiability. By understanding the challenges posed by non-identifiability and employing appropriate methods, researchers can improve the reliability and interpretation of their disease models.
An alternative approach to addressing non-identifiability is to reframe the problem by focusing on identifiable parameters or combinations of parameters. For instance, Gustafson [6] outlines methods for partially identified models, where inference is restricted to identifiable quantities. Additionally, marginalising over non-identifiable parameters within a Bayesian framework can still provide useful insights about identifiable components. These strategies complement the use of informative priors and may be particularly useful in cases where eliciting prior information is challenging.
To address the challenges of non-identifiability, this tutorial focuses on incorporating informative prior distributions for model parameters. This approach is supported by the work of Neath and Samaniego [38], who demonstrated the effectiveness of prior information in handling a non-identifiable model. They have demonstrated that the extent of the improvement that a posterior estimate provides over a prior estimate is potentially quite large. However, the potential for the posterior estimate of a non-identifiable parameter being inferior to the prior because of Bayesian updating is bounded. By specifying appropriate priors, we can constrain the parameter space and improve the precision of our posterior estimates.
3 Incorporating Prior Information to Address Non-Identifiability
@@
In practice, non-identifiability can be diagnosed when there is a lack of convergence of chains produced by the Metropolis-Hastings algorithm. For example, when the posterior distributions of parameters are flat and broad, or when there is cross-correlation between sampled values.
To assess the convergence of MCMC chains, visual diagnostics of trace plots such as Figure 2 and Figure 5 are crucial. Trace plots provide valuable insights into the chains’ behaviour. They help visualise the evolution of parameter values over iterations, with convergence indicated by a stable, random walk. A well-converged chain will exhibit a stable, random walk with no discernible trends (the right-hand side of both plots). Conversely, trends, cycles, or abrupt jumps indicate non-convergence (the left-hand side of both plots). By examining these plots, we can identify potential issues like non-convergence or poor mixing and take corrective actions.
@@
To resolve non-identifiability, the suggested approach is to include information about some of the variables where this is obtainable. This can be done by carefully considering whether additional information has been collected or is available from other sources. For example, duration of illness may not be available locally, but a prior may be obtained by considering international evidence that is emerging. Note that from a structural and inferential viewpoint, the prior contains information about the parameters that is not available in the likelihood. It does not necessarily have to have arisen or been obtained before the data which is used in the likelihood function. As in the examples above, observing cases directly and noting the time to recovery will give us different information than prevalence estimates when trying to ascertain rates of incidence or occurrence of disease.
In algorithmic terms, the process of calibration would proceed as follows:
@@
  • Identify model outputs of interest (targets) and populate these with collected data.
@@
  • Run the MCMC calibration process.
@@
By effectively incorporating prior information, we can improve the precision and reliability of parameter estimates in SIS models, especially when dealing with non-identifiability issues.
@@ 4 Example: SIS model
To illustrate the challenges of non-identifiability further we provide an example in line with the model introduced in Figure 1(a). We consider a simple susceptible-infected-susceptible (SIS) compartmental model. This simple SIS compartmental model represents the dynamics of diseases with no permanent immunity. In this model, individuals can move between the susceptible (S) and infected (I) compartments. The transitions between the two states are described by
@@
where $\beta=c\cdot p$ and $\gamma=1/d$ with $c$ being the contact rate, $p$ the transition probability and $d$ the infectious period. Most of the population starts off in the susceptible state with $S=0.99$ and $I=0.01$. We treat this situation as a population where there has initially been disease introduced, and at some later stage, a stable situation has been observed where 15,000 of 25,000 individuals are identified as infected. The likelihood may be summarised as probabilities of being susceptible or infected (see Figure 1(c) for an example of disease dynamic). After a period of active disease dynamics, the probabilities are stabilised around the 0.4 and 0.6 values for infected and susceptible compartments respectively. From the data collection point (i.e. observing 15,000 infected individuals in a population of 25,000 with some uncertainty added for observational error), this could be transformed into having normal distributions corresponding to probabilities of being in each compartment introduced as targets. These targets ($N(0.4,0.01)$ and $N(0.6,0.01)$) are then used in the Metropolis-Hastings algorithm. In particular, this means that the full SIS model is run for every proposed new input parameter combination and the produced output is compared to the target distributions. The likelihood is calculated as the negative log of the product of the two probability density functions of the two values output by the model normalised by the values of those density functions at the modes. Implicit in this is that a uniform prior is used for $p$ and $c$.
The SIS model can exhibit non-identifiability, particularly when the data are limited or noisy. For example, it can be difficult to distinguish between high transmission rates with high recovery rates and low transmission rates with low recovery rates if the overall prevalence of the disease remains relatively constant as described in Section 2.4. To address non-identifiability in models, prior information can be invaluable. This information can come from various sources, including:
  • Expert knowledge. Epidemiologists or infectious disease experts may have prior beliefs about the transmission rate ($\beta$) based on their understanding of the disease, its biology, and previous outbreaks.
  • Historical data. Data from past outbreaks of similar diseases can provide insights into the likely range of transmission rates.
  • Biological constraints. Biological factors, such as the average contact rate or the duration of infectiousness, can be used to inform the prior distribution for $\beta$.
To incorporate prior information into the SIS model, we can specify a prior distribution for the transmission rate ($\beta$). For example, if expert opinion suggests that $\beta$ is likely to be between 0.4 and 0.6, we could use a uniform distribution over this range.
It is essential to assess the impact of the prior distribution on the posterior estimate of $\beta$. Sensitivity analysis can be used to evaluate how the posterior varies with different prior specifications. If the posterior is highly sensitive to the prior, it suggests that the data are not very informative and that the prior is playing a dominant role in shaping the estimate. That means that without a strong informative prior it will be impossible to have a focused posterior and therefore make any meaningful inference.
@@
After running a simple Metropolis-Hastings algorithm for $c,\ p$ and $d$ with a uniform prior and the two normal distributions ($N(0.4,0.01)$ and $N(0.6,0.01)$) as our targets we plot the resulting $\beta$ and $\gamma$ values in Figure 1(b). It is noticeable that those are almost perfectly correlated. In panel (b) of the figure, the impact of specifying a prior for one variable is shown. The density in panel (b) represents an informative prior, with arrows illustrating how it impacts upon the joint posterior and hence the marginal posterior of $\gamma$. The annotated density on the horizontal axis for $\beta$ illustrates how the information in that parameter restricts the joint posterior, and thus provides indirect information on $\gamma$, the marginal distribution of which is now affected. Even though no information about the second variable is added, conditioning on a narrower prior distribution for $\beta$ restricts the joint posterior space to a narrower range, thus impacting the marginal posterior distribution for $\gamma$ which is induced due to the correlation between the parameters. Additionally, because $\beta=c\cdot p$, $c$ and $p$ are perfectly correlated. This leads to the chains for $c,\ p$ and $d$ not converging due to the complex links between their values and the observational equivalence of the resulting targets.
Figure 1: Simple example of an SIS model (a). Panel (b) shows the correlation between two parameters and the effect of a prior for one parameter ($\beta$) on the joint posterior space and marginal posterior distribution of $\gamma$. Panel (c) presents disease dynamics over time.
@@
Figure 2: Resulting chains of running the Metropolis-Hastings algorithm with improper uniform priors and with informative priors for the SIS model.